Loan Amount:               $10,000,000.00

Interest Rate:                6-3/4%

Amortization:               30 years

Term:                           10 years

  1. Assume the above loan is interest only for the first three (3) years.
    1. What is the Annual Debt Service (ADS) in the first year?
    2. How much interest (cumulative) will have been paid at the end of year 3?
    3. How much principal will be due at maturity?
    4. How much interest will have been paid over the 10 year term?
Answer #1

interest rate per month = 6.75% / 12 = 0.5625

for 36 months i.e 3 years = 10,000,000 * 0.5625 = 56250

a) annual debt service in 1st year = 56250*12 = 675000

b) cumulative interest pain at end of 3rd year= 675000*3 = 2025000

c) principal due at maturity = 8849998.57 (value from amortization chart )

d) interest will be paid over 10 year term = 6499064.79 ( value from amortization chart )